OpenStudy (loser66):

Can I write (j+1)(j+2)(j+3)......(j+n) = (j+n)! Please

OpenStudy (kc_kennylau):

Only if j=0 or j=1

OpenStudy (kc_kennylau):

P.S. (j+1)(j+2)(j+3)......(j+n) = (j+n)! / j!

OpenStudy (loser66):

Thank you

OpenStudy (kc_kennylau):

no problem

OpenStudy (loser66):

I have problem with express the stuff: S, D are linear transformations on P P(t) = \(\sum_{j=0}^{n-1} \xi_j t^j\) SP(t) = \(\sum_{j=0}^{n-1} \dfrac{\xi_j}{j+1} t^{j+1}\) DP(t) = \(\sum_{j=0}^{n-1} j\xi_j t^{j-1}\) calculate \(D^nS^n\) and \(S^nD^n\)

OpenStudy (loser66):

If using D, S as differentiation and integration, it's quite easy to calculate. In this case, I have to use those notations to calculate :(

OpenStudy (loser66):